OpenStudy (dls):

Integrate log(1+cosx)dx from 0 to pi

OpenStudy (dls):

\[\Huge \int\limits_{0}^{\pi} \log(1+cosx)\]

OpenStudy (sumi29):

This requires a bit of trick. Hint: \[\int\limits_{a}^{b}\log_{n}x = (x \log_{n}x-x/\ln(n))+C \]

OpenStudy (dls):

I am at \[\Huge I=\int\limits_{0}^{\pi}logsinx\]

OpenStudy (dls):

dunno what to do next

OpenStudy (sumi29):

Well do you know how to integrate by parts? If you don't just post here and I will help you solve this step by step.

OpenStudy (experimentx):

\[ 1 + \cos (x) = 2 \cos^2(x/2)\] make use of log function and change of variables. If you know how to integrate log(sin(x)) from 0 to pi then you should know this too.

OpenStudy (experimentx):

\[ I = \int_0^\pi \log(\sin(x))dx = \log(2) \pi + \int_0^\pi \log(\cos(x/2))dx + \int_0^\pi \log(\sin(x/2))dx \\ = \pi \log(2) + 2 \int_0^{\pi/2} \log(\cos(x))dx + 2 \int_0^{\pi/2} \log(\sin(x))dx \\ = \pi \log(2) + 2 \left (\int_0^{\pi/2} \log(\cos(x))dx + \int_0^{\pi/2} \log(\sin(x))dx \right ) \\= \pi \log(2) + 2 \left (\int_0^{\pi/2} \log(\sin(x+\pi/2))dx + \int_0^{\pi/2} \log(\sin(x))dx \right ) \\=\pi \log(2) + 2 \left (\int_{\pi/2}^\pi \log(\sin(x))dx + \int_0^{\pi/2} \log(\sin(x))dx \right ) \\=\pi \log(2) + 2I \]